OpenStudy (anonymous):

A capacitor has a plate area of 0.01 square meters. A second capacitor has a plate area of 0.02 square meters. If the first capacitor is 0.1 micro Farads, what is the value of the second capacitor if all other physical characteristics are identical?

OpenStudy (sidsiddhartha):

use this \[capacitence,C=\frac{ \epsilon_0*A }{ d }\\A=area,d=distence~ Between~plates\]

OpenStudy (sidsiddhartha):

so capacitence is directly proportional to the area

OpenStudy (anonymous):

I haven't been given d in the question though

OpenStudy (sidsiddhartha):

\[\frac{ C_1 }{ C_2 }=\frac{A_1}{A_2}\] now ,everything is given C_1=0.1, A1=0.01 A2=0.02 just plug them in and find C2

OpenStudy (anonymous):

thank you!

OpenStudy (sidsiddhartha):

yw!!